A contractor employed a certain number of men to complete a work in a certain number of days.If three more men were deployed the work would be finished two days earlier while if twelve more men were deployed it would be finished five days earlier.Find number of men deployed by the contractor

Grade:10

1 Answers

Sunil Raikwar
10 years ago
Use the formula m1d1=m2d2 in such types of problems where m means no. of workers & d means no of days.
Let no. of man be x & no. of days be y, then
xy=(x+3)(y-2)
3y-2x=6........1
& xy=(x+12)(y-5)
12y-5x=60............2
Solving eq 1&2 we get x=12& y=10
Thanks and Regards
Sunil Raikwar
askIitians faculty